OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (–2, 1) and (4, 13)? y – 1 = –2(x – 4) y – 1 = 2(x + 2) y – 13 = –2(x + 2) y – 13 = 2(x + 4) 2.What is the equation of the line passing through (–3, 6) and parallel to the line y = 4x – 2 in standard form? x – 4y = –18 x + 4y = 18 4x – y = –18 4x + y = 18 PLEASE HELP IM IN A HURRY I'LL GIVE A MEDAL

OpenStudy (ivettef365):

I already gave you answer for the first one, did you get it??

OpenStudy (anonymous):

FInd the slope between the two points first: m = y2-y1/x2-x1 and plug that into y-y1=m(x-x1)

OpenStudy (anonymous):

No i didn't ivette

OpenStudy (anonymous):

y – 1 = 2(x + 2) ? right

OpenStudy (ivettef365):

correct

OpenStudy (ivettef365):

for the second equation what is the slope on equation y=4x-2

OpenStudy (anonymous):

the slop of y=4x-2 is 4/1